### v1

# Improving Discrete Diffusion Unmasking Policies Beyond Explicit Reference Policies

Abstract
Masked diffusion models (MDMs) have recently emerged as a novel framework for language modeling. MDMs generate sentences by iteratively denoising masked sequences, filling in [MASK] tokens step by step. Although MDMs support any-order sampling, performance is highly sensitive to the choice of which position to unmask next. Prior work typically relies on rule-based schedules (e.g., max-confidence, max-margin), which provide ad hoc improvements. In contrast, we replace these heuristics with a learned scheduler. Specifically, we cast denoising as a KL–regularized Markov decision process (MDP) with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. We prove that the optimized policy under this framework generates samples that more closely match the data distribution than heuristic schedules. Empirically, across four benchmarks, our learned policy consistently outperforms max-confidence: for example, on Sudoku, where unmasking order is critical, it yields a 20.1% gain over random and a 11.2% gain over max-confidence.
1 Introduction
Diffusion models have achieved remarkable success in domains such as image generation (Ho et al., 2020; Song et al., 2021), by defining a noise-adding forward process and a denoising reverse process in continuous space, and learning the reverse dynamics within this framework. Recent studies (Nie et al., 2025; Ye et al., 2025; Lou et al., 2024) have advanced this line of work into the discrete domain, applying it to language modeling through masked diffusion models (MDMs). In the discrete setting, MDMs define a forward process that replaces tokens in text with a [MASK] token and a corresponding reverse process; training typically maximizes an evidence lower bound (ELBO) (Lou et al., 2024) to fit the reverse dynamics. Small-scale MDMs (Lou et al., 2024; Sahoo et al., 2024) have matched or surpassed autoregressive models (ARMs), and recent large-scale studies (Ye et al., 2025; Nie et al., 2025) report results competitive with strong ARMs such as LLaMA-3 (Grattafiori et al., 2024).
Despite a shared template, diffusion in discrete versus continuous spaces differs substantially—especially at inference. Continuous-space diffusion models (Song et al., 2021; Karras et al., 2022) perform denoising by integrating a stochastic differential equation (SDE) or an ordinary differential equation (ODE). In contrast, discrete-space diffusion models (Sahoo et al., 2024; Nie et al., 2025) denoise by iteratively sampling concrete replacements for [MASK] tokens. These mechanisms lead to different update geometries: continuous models accumulate numerous small changes across all dimensions, whereas discrete models enact localized, high-magnitude jumps at a few coordinates. Consequently, effective denoising in discrete space requires its own methodology. Just as advances in SDE solvers unlocked large gains for continuous diffusion, improved denoising strategies are likely to yield substantial performance gains for MDMs.
This raises a natural question: what matters most in MDM denoising? Among several factors, we focus on the choice of which position to unmask next. Kim et al. (2025) provide a theoretical rationale for this choice, proving that no polynomial-time algorithm can solve any-order generation; i.e., we cannot train MDMs that exactly recover the real data distribution for all masked sentences. However, they empirically show that selecting unmasking tokens via max-confidence (Chang et al., 2022) or max-margin (Kim et al., 2025) at inference time can bypass sub-hard instances and even outperform ARMs. Reflecting this, contemporary large-scale MDMs such as LLaDA or Dream-7B (Nie et al., 2025; Ye et al., 2025) commonly adopt max-confidence.
In this paper, we aim to move beyond heuristic references (e.g., max-confidence) by learning improved unmasking policies. To this end, we cast MDM denoising as a Markov decision process (MDP) and train an unmasking position-selection policy using the group relative policy optimization (GRPO) (Shao et al., 2024). Theoretically, we prove that under this framework, the optimized policy produces samples closer to the true data distribution than the strong heuristic references. Moreover, we introduce three tractable surrogate objectives and provide propositions that justify optimizing them in place of the ideal, yet intractable, objective. Practically, our learned policy substitutes for max-confidence and related heuristics, achieving improved accuracy across four benchmarks. For example, on Sudoku, where unmasking order is crucial, it delivers a 20.1% improvement over random and a 11.2% improvement over max-confidence.
Related works. Reinforcement learning (RL) has emerged as an effective tool to fine-tune large language models toward complex objectives or preferences (Ouyang et al., 2022; Guo et al., 2025; Shao et al., 2024; Rafailov et al., 2024). In MDMs, Zekri & Boullé (2025) apply policy gradient methods to a pretrained discrete diffusion model so as to maximize a task-specific reward at the last denoising step. Zhao et al. (2025) adopts the GRPO framework to improve the reasoning ability by rewarding only the final answer accuracy. Zhu et al. (2025) propose a variance reduced preference optimization (VRPO) for MDMs, which shares the concept of direct preference optimization (DPO) (Rafailov et al., 2024) in ARMs. While the aforementioned works utilize RL to post-train the MDM itself, they differ from ours, where we train the unmasking policy model. Recently, Huang et al. (2025) introduce DColT, which also trains an unmasking policy via reinforcement learning. Our approach differs in that we formulate denoising as a KL–regularized MDP with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. Furthermore, while Huang et al. (2025) jointly applies RL to train both the unmasking policy and the underlying MDMs for maximum accuracy, our work focuses solely on optimizing the unmasking policy for frozen MDMs, thereby reducing computational cost and mitigating the risk of reward over-optimization.
2 Preliminary
Notations. Let the distribution $p_{\rm{data}}$ be the data distribution over sequences of length $L$ with vocabulary $\mathcal{V}=\{1,...,m\}$, and let the set of all possible sequences be $\mathcal{X}=\mathcal{V}^{L}$. We define $p_{L}$ as the Kronecker delta distribution that assigns all probability mass to the sentence in which every token is masked. We denote the kernel forms of $p_{\rm{data}}$ and $p_{L}$ by ${\mathbf{P}}_{\rm{data}}$ and ${\mathbf{P}}_{L}$, respectively, with entries ${\mathbf{P}}_{\rm{data}}]{{\bm{x}}}=p_{\rm{data}}({\bm{x}})$. We write ${\bm{x}}_{n}=\{x_{n}^{1},x_{n}^{2},\dots,x_{n}^{L}\}$ for the $n$-th step sequence with $L$ tokens. We denote the mask by $\mathbb{M}$ and the set of sequences with $n$ masks by $\mathcal{X}_{n}=\{{\bm{x}}|{\bm{x}}\in\{\mathcal{V}\cup\{\mathbb{M}\}\}^{L},\sum_{i=1}^{L}{\textbf{1}(x^{i}=\mathbb{M})}=n\}$, so that ${\bm{x}}_{n}\in\mathcal{X}_{n}$ has $n$ masks and we refer $a^{i}[{\bm{x}}]$ for the $i$-th mask index of $\bm{x}$}. For example, if the number of masks in $\bm{x}$} is $n$, then $a^{i}[{\bm{x}}]\in\{1,\dots,L\}$, $x^{a^{i}[{\bm{x}}]}=\mathbb{M}$ and $i\in\{1,\dots,n\}$. For brevity, we will omit $\bm{x}$} in $a^{i}[{\bm{x}}]$ for the rest of the paper. We denote MDMs as $\theta$ and their model distribution as $\pi_{\theta}$. We denote Markov transition kernel ${\mathbf{T}}:\mathcal{X}\times\mathcal{X}\to[0,1]$ where ${\mathbf{T}}({\bm{x}},{\bm{x}}^{\prime})$ is the probability of $\bm{x}$} transform into $\bm{x}$$^{\prime}$ such that $\sum{\mathbf{T}}_{{\bm{x}}^{\prime}\in\mathcal{X}}({\bm{x}},{\bm{x}}^{\prime})=1$ for all $\bm{x}$}.
Masked diffusion models. Masked Diffusion Models (Shi et al., 2024; Sahoo et al., 2024) define a mask-adding forward process and corresponding reverse process in discrete space, and learn the model $\pi_{\theta}$ to mimic the true reverse process. The forward process gradually masks tokens independently one-by-one in ${\bm{x}}_{0}$ until the sequence is fully masked at ${\bm{x}}_{n}$ with $n=L$ (or $t=1$ where $t:=n/L$). For $t=n/L\in(0,1)$, the sequence ${\bm{x}}_{n}$ is partially masked, with each being masked with probability $t$ or remaining unmasked with probability 1t1-t. The reverse process recovers the data distribution by iteratively predicting masked tokens as $t$ decreases from $1$ to 0. While previous works (Campbell et al., 2022; Sahoo et al., 2024) propose the time-dependent MDM training framework, the recent works (Zheng et al., 2025; Ou et al., 2025) propose the time-agnostic model where time is implicitly included in the number of masks. Then, the time-agnostic loss for MDM is defined as follows:
$\textstyle\mathcal{L}_{\rm{MDM}}(\theta)\triangleq-\mathbb{E}_{n,{\bm{x}}_{0},{\bm{x}}_{n}}\big[\tfrac{1}{n}\sum_{i=1}^{L}\textbf{1}[x_{n}^{i}=\mathbb{M}]\log\pi_{\theta}(x_{0}^{i}|{\bm{x}}_{n})\big].$ (1)
Zheng et al. (2025); Ou et al. (2025) have shown that $\mathcal{L}_{\rm{MDM}}$ is the upper bounds of the negative log-likelihood. Zheng et al. (2025) propose the simple generation algorithm, named the First-Hitting sampler, beyond a continuous time Markov chain (CTMC) that follows the conventional reverse process. First-Hitting sampler randomly selects the index $i$ for unmasking and samples the next token sequentially as ${\bm{x}}_{L}\rightarrow{\bm{x}}_{L-1}\rightarrow\dots\rightarrow{\bm{x}}_{0}$. This is proven to converge to $p_{\rm{data}}$. Recent large-scale MDMs (Nie et al., 2025; Ye et al., 2025) currently adopt time-agnostic modeling and further employ sampling strategies such as the max-confidence beyond the First-Hitting sampler.
Markov transition kernel of random sampler and training hard problem of MDM. Large-scale MDMs first select the masked token to unmask and sample from the categorical distribution. See Algorithm 1 in Appendix B.2 for the detailed sampling algorithm. We here formulate general Markov transition kernels defined by various samplers of MDM.
${\mathbf{T}}_{n}^{(g)}=\sum\limits_{i=1}^{n}{{\mathbf{T}}_{n}^{i,(g)}},\;\;\;\;{\bm{x}}_{n-1}\sim\sum\limits_{i=1}^{n}[{\mathbf{T}}_{n}^{i,(g)}]({\bm{x}}_{n},{\bm{x}})$ (2)
where
$[{\mathbf{T}}_{n}^{i,(g)}]({\bm{x}}_{n},{\bm{x}})=\left\{\begin{array}[]{cl}g(a^{i}|{\bm{x}}_{n})\cdot\pi_{\theta}(x^{a^{i}}|{\bm{x}}_{n})&\text{if }{\bm{x}}_{n}^{-{a^{i}}}={\bm{x}}^{-a^{i}},\\ 0&\text{otherwise},\end{array}\right$. (5)
where $\sum_{i=1}^{n}g(a^{i}|{\bm{x}}_{n})=1$, ${\bm{x}}^{-{a^{i}}}=\{x^{1},x^{2},\dots,x^{{a^{i}}-1},x^{{a^{i}}+1},\dots,x^{L}\}$, $x_{n}^{a^{i}}=\mathbb{M}$ by definition, and $\pi_{\theta}(x_{n-1}^{a^{i}}|{\bm{x}}_{n})$ is the categorical distribution of model prediction of the $i$-th mask of ${\bm{x}}_{n}$.
The non-homogeneous Markov kernel for First-Hitting sampler (Zheng et al., 2025) for time-agnostic MDMs can be defined as choosing $g$ by
$g(a^{i}|{\bm{x}}_{n}):=g_{\text{rand}}(a^{i}|{\bm{x}}_{n})=\frac{1}{n}$ (7)
In algorithmic manner, this is equal to first sample $i\sim\mathcal{U}\{1,\dots,n\}$, then sample $x^{a^{i}}\sim\pi_{\theta}(x^{a^{i}}|{\bm{x}}_{n})$, and obtain ${\bm{x}}_{n-1}$ by replacing $i$-th mask of ${\bm{x}}_{n}$ with $x^{a^{i}}$. In this case, $(\prod_{n=1}^{L}{\mathbf{T}}_{n}^{g_{\rm{rand}}})\cdot{\mathbf{P}}_{L}\approx{\mathbf{P}}_{\rm{data}}$ if $\pi_{\theta}$ is perfectly estimated  (Zheng et al., 2025). However, Kim et al. (2025) have proven that learning $\pi_{\theta}$ for every combination of masked sequence cannot be solved by a polynomial algorithm, and there exist hard subproblems. This can interpreted as there exists some points that true reverse process and model’s reverse process are different, such that resulting in $(\prod_{n=1}^{L}{\mathbf{T}}_{n}^{g_{\rm{rand}}})\cdot{\mathbf{P}}_{L}\neq{\mathbf{P}}_{\rm{data}}$. We provide a simplified example to help understand:
Example 1.
Consider the following simplified zebra (Einstein) puzzle. There are two houses, and the task is to determine the person living in each house and their favorite food. The possible persons are Robert and Tom, and the possible foods are pizza and hamburger. The clues are given as follows: 1) The person living in the first house is Robert, and 2) Tom likes pizza.
Now consider the MDM directly answer the questions, filling the mask of sentence: “House 1: name: [MASK], food: [MASK], House 2: name: [MASK], food: [MASK]”. Obviously, if MDM tries to answer the name of the person who lives in the first house, the problem becomes trivial. However, if it tries to fill attributes for the second house first before filling attributes of the first house, the problem becomes much harder. As the number of questions and clues grows, this kind of discrepancy would be much larger. Indeed, the zebra puzzle is also known as an NP-hard problem (Shah et al., 2024; Yato & Seta, 2003). Note that this is a highly simplified example to help understand the problem of MDM training, and refer to Kim et al. (2025) for a complete explanation.
Practical approach to avoid such hard problems. Meanwhile, Kim et al. (2025) empirically shows that unmasking the token with max-confidence or max-margin helps to avoid such hard problems, leading to better generations. Therefore, in practice, rather than randomly sampling the token for unmasking, large-scale discrete diffusion models (Nie et al., 2025; Ye et al., 2025) utilize such rule-based samplings. The weighted kernel related to the max-confidence sampling can be defined with $g(a^{i}|{\bm{x}}_{n}):=g_{\text{conf}}$ defined as follows:
$g_{\text{conf}}(a^{i}|{\bm{x}}_{n})=\textbf{1}\bigl(a^{i}=\operatorname*{arg\,max}_{a^{j}}(\max_{c\in\mathcal{V}}\pi_{\theta}({\bm{x}}_{n-1}^{a^{j}}=c|{\bm{x}}_{n}))\bigr),$ (8)
or in other stochastic forms,
$g_{\rm{conf}^{\tau}}(a^{i}|{\bm{x}}_{n})$ $=\frac{\sum_{c\in\mathcal{V}}\exp({\pi_{\theta}(x_{n-1}^{a^{i}}=c|{\bm{x}}_{n})/\tau})}{\sum_{a^{j}}\sum_{c\in\mathcal{V}}\exp({\pi_{\theta}(x_{n-1}^{a^{j}}=c|{\bm{x}}_{n})/\tau})},\;\;\text{s.t.}\;\lim_{\tau\rightarrow 0+}g_{\rm{conf}^{\tau}}=g_{\text{conf}},$ (9)
$g_{\rm{Top-K}}(a^{i}|{\bm{x}}_{n})$ $=\frac{1}{K}\cdot\textbf{1}(a^{i}\in\underset{{a^{j}}}{\rm{argtopk}}(\max_{c\in\mathcal{V}}\pi_{\theta}({\bm{x}}_{n-1}^{a^{j}}=c|{\bm{x}}_{n}))),\;\;\text{s.t.}\;g_{\rm{Top-1}}=g_{\rm{conf}}.$ (10)
where $\rm{argtopk}$ denotes the indices from the top-k sampler. Similarly, the weight kernel for max-margin sampler can be defined with $g_{\text{margin}}$ whose value is $1$ where $i$ is the mask index with the maximum margin between the first-confident and second-confident token.
3 Improving Unmasking Policy Beyond a Reference Policy
While twisted Markov transition kernels driven by heuristic unmasking policies—such as $g_{\rm conf}$ or $g_{\rm margin}$—deliver substantial gains, we posit that a stronger policy $g$ can guide MDM denoising more effectively. Motivated by this, we learn $g_{\phi}$ instead of defining $g$ in a heuristic manner to find an optimal path. We further cast the learning problem of $g_{\phi}$ as a KL-regularized MDP, where a strong heuristic sampler serves as a reference policy, enabling stability and policy improvement guarantees.
3.1 Existence of optimal path Beyond $g_{\rm{conf}}$
Figure 1: Pass@N on GSM8K. The dashed red line is the single-trajectory accuracy of max-confidence $g_{\rm conf}$.
We test various $g$ in GSM8K (Cobbe et al., 2021), the mathematical questions benchmark, to probe whether max-confidence is merely strong or whether substantially better unmasking paths exist. We compare three schedulers: $g_{\rm conf}$, $g_{\rm rand}$, and $g_{\rm{Top-K}}$. Here, $g_{\rm conf}$ is deterministic, while $g_{\rm Top\text{-}K}$ smoothly interpolates between random and max-confidence.
For each test question, we draw $N$ independent trajectories using either $g_{\rm rand}$ or $g_{\rm Top\text{-}5}$ and report Pass@N, the fraction of questions for which at least one trajectory is correct. Figure 1 shows three consistent trends: (i) $g_{\rm{conf}}$ significantly outperforms $g_{\rm{rand}}$ when $N=1$; (ii) $g_{\rm Top\text{-}5}$ dominates $g_{\rm rand}$ across $N$, highlighting the strength of confidence-based ordering; (iii) as $N$ grows, both curves surpass the max-confidence line by a clear margin, and $g_{\rm Top\text{-}5}$ exceeds $0.92$ when $N=10$.
These results carry two implications. First, max-confidence is indeed a strong single-path baseline. Second, the rise of Pass@N above the max-confidence line indicates the presence of higher-yield unmasking trajectories than the deterministic max-confidence order. Brute-force search for such paths is computationally prohibitive and requires unobserved oracle feedback, motivating our learned policy $g_{\phi}$ that aims to discover high-reward paths without enumerating many trajectories.
3.2 Reformulating the problem as reinforcement learning
Motivated by the strength of $g_{\rm conf}$ and the existence of higher-yield unmasking paths, we set two goals: (i) increase the probability that an MDM produces the correct answer, and (ii) move the induced sampling distribution closer to $p_{\rm data}$ than a strong heuristic reference $g_{\rm ref}$. We therefore learn a stochastic position-selection policy $g_{\phi}$ rather than fixing $g$ heuristically. The policy is parameterized as a softmax over masked positions using a learnable scorer $h_{\phi}({\bm{x}})\in\mathbb{R}^{L}$:
$g_{\phi}(a^{i}\,|\,{\bm{x}}_{n})=\frac{\exp\!\big([h_{\phi}({\bm{x}}_{n})]_{a^{i}}\big)}{\sum_{i^{\prime}=1}^{n}\exp\!\big([h_{\phi}({\bm{x}}_{n})]_{a^{i^{\prime}}}\big)}.$ (11)
We cast unmasking as a regularized Markov decision process (MDP) with verifiable terminal rewards. Each episode begins from a prompt–answer pair where the answer is fully masked, ${\bm{x}}_{L}:{\mathbf{q}}]$ with ${\mathbf{q}}\sim\rho_{\mathcal{Q}}$ referring to the answer distribution and ${\bm{x}}_{L}=\mathbb{M}^{L}$, and terminates when all masks are removed at $n=0$, yielding ${\bm{x}}_{0}\in\mathcal{X}=\mathcal{V}^{L}$. At state ${\bm{x}}_{n}$ the action space is the set of masked indices $\mathcal{A}_{{\bm{x}}_{n}}=\{a^{1}[{\bm{x}}_{n}],\dots,a^{n}[{\bm{x}}_{n}]\}$; the policy $g_{\phi}(\cdot\,|\,{\bm{x}}_{n})$ selects one index $a_{n}\in\mathcal{A}_{{\bm{x}}_{n}}$ to unmask. Transitions factor through the fixed MDM denoiser $\pi_{\theta}$:
$p_{g_{\phi}}({\bm{x}}_{n-1}\mid{\bm{x}}_{n})={\mathbf{T}}_{n}^{g_{\phi}}({\bm{x}}_{n},{\bm{x}}_{n-1})=g_{\phi}(a_{n}\mid{\bm{x}}_{n})\cdot\pi_{\theta}({\bm{x}}_{n-1}\mid{\bm{x}}_{n},a_{n}),$
where $\pi_{\theta}({\bm{x}}_{n-1}\mid{\bm{x}}_{n},a_{n})$ is the token distribution for position $a_{n}$. The agent thus controls where to unmask, while the environment dynamics are induced by $\pi_{\theta}$. Rewards are verifiable: an external checker evaluates the fully unmasked output ${\bm{x}}_{0}$ and returns $r({\mathbf{q}},{\bm{x}}_{0})\in\{0,1\}$.
To stabilize and accelerate training, we introduce a strong reference policy $g_{\rm ref}$ (e.g., Top-$K$) and optimize a theoretical KL-regularized GRPO-style objective (Mroueh, 2025). We optimize directly over the terminal-output distribution induced by the policy: all expectations are taken with respect to $p_{g}({\bm{x}}_{0}\mid{\mathbf{q}})$, and the regularizer compares terminal-output marginals. Accordingly, our theoretical loss is an output-level GRPO objective without clipping:
$\max_{\phi}\;\;\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\Biggl[\mathbb{E}_{{\bm{x}}_{0}\sim p_{g_{\phi_{\rm old}}}(\cdot\mid{\mathbf{q}})}\!\biggl[\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0}\mid{\mathbf{q}})}\,A({\mathbf{q}},{\bm{x}}_{0})\biggr]-\beta\,D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\big\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)\Biggr],$ (12)
where $A({\mathbf{q}},{\bm{x}}_{0})$ is a standardized advantage. Let $r_{g}({\mathbf{q}})=\mathbb{E}_{{\bm{x}}_{0}\sim p_{g}(\cdot\mid{\mathbf{q}})}[r({\mathbf{q}},{\bm{x}}_{0})]\in[0,1]$ and $\mathrm{std}_{g}({\mathbf{q}})=\sqrt{\mathbb{E}_{{\bm{x}}_{0}\sim p_{g}(\cdot\mid{\mathbf{q}})}[(r({\bm{x}}_{0})-r_{g}({\mathbf{q}}))^{2}]}$, then $A({\mathbf{q}},{\bm{x}}_{0})=(r({\mathbf{q}},{\bm{x}}_{0})-r_{g_{\phi_{\rm old}}}({\mathbf{q}}))/(\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon)$. This objective follows the output-level GRPO loss used in DeepSeek-R1 (Guo et al., 2025) rather than the original token-wise GRPO used in DeepSeekMath (Shao et al., 2024). The regularizer keeps $g_{\phi}$ close to $g_{\rm ref}$, acting as a trust-region that mitigates instability while enabling improvement over a strong baseline; selecting an appropriate $g_{\rm ref}$ provides a high-accuracy starting point and accelerates convergence.
3.3 Theoretical analysis
In this section, we provide how $g_{\phi}$ optimizing Eq. (12) can guarantee the convergence of the policy model and superior performance over the reference policy $g_{\rm{ref}}$. Mroueh (2025) have proven that a policy model learning output-level GRPO loss (Eq. (12)) can converge to its optimal performance $r_{g^{*}}$, which is higher than $r_{g_{\rm{ref}}}$:
Theorem 1 (Restatement of GRPO convergence theorem (Mroueh, 2025)).
Assume $1>r_{g_{\rm{ref}}}>0$, define for $\beta>0$:
$h_{\epsilon,p_{\rm{ref}}}(r_{g})=\frac{1}{1+\frac{1-r_{g_{\rm{ref}}}}{r_{g_{\rm{ref}}}}\exp\bigr(-\frac{1}{\beta}\frac{1}{\sqrt{r_{g}(1-r_{g})+\epsilon}}\bigl)}$ (13)
[…]
$r_{g_{\phi_{n}}}({\mathbf{q}})=h_{\epsilon,p_{\rm{ref}}}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))$ (14)
and $r_{g_{\phi_{0}}}({\mathbf{q}})=r_{g_{\rm{ref}}}({\mathbf{q}})$. Also, let $r_{g^{*}}$ be a fixed point of $h_{\epsilon,p_{\rm{ref}}}$ and assume that have $|h_{\epsilon,p_{\rm{ref}}}^{\prime}(r_{g^{*}})|<1$. Given that $h_{\epsilon,p_{\rm{ref}}}$ and $h_{\epsilon,p_{\rm{ref}}}^{\prime}$ are continuous in $[0,1]$, following local fixed point convergence is established:
$\lim_{n\rightarrow{\infty}}r_{g_{\phi_{n}}}=r_{g^{*}},$ (15)
and $r_{g^{*}}>r_{g_{\rm{ref}}}$.
Therefore, setting $g_{\rm{ref}}$ as a strong proxy model is important, as we can always get a better policy $g_{\phi^{*}}$. Moreover, in the following we prove that policy model $\phi$ maximizing Eq. (12) can restore $p_{\rm{data}}$ better than $g_{\rm{ref}}$:
Theorem 2 (Reference-KL Tightening for MDM Policy Improvement).
Assume there exists ideal MDM $\theta^{*}$ satisfying $\mathcal{L}_{\rm{MDM}}(\theta^{*})=0$ for $p_{\rm{data}}$ where data is composed of prompt and oracle answer $\{{\mathbf{q}},{\mathbf{a}}\}$ such that $r({\mathbf{q}},{\mathbf{a}})=1$. Now consider incomplete MDM $\theta$ such that $\mathcal{L}_{\rm{MDM}}(\theta)>0$, and $g_{\phi^{*}}$ that satisfies Eq. (15) for $\theta$. Let let $g_{\mathrm{ref}}$ be a policy model such that, for every $\mathbf{q}$}, $p_{g_{\rm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$ whenever $p_{\theta^{*}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$, i.e., $\operatorname{supp}(p_{\theta^{*}}(\cdot\mid{\mathbf{q}}))\subseteq\operatorname{supp}(p_{g_{\mathrm{ref}}}(\cdot\mid{\mathbf{q}}))$. Then, the following inequality holds:
$D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))\leq D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (16)
Or equivalently,
$D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))\leq D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (17)
Proof sketch..
Mroueh (2025) provides policy dynamic under training Eq. (12). Using the result given in Lemma 1 in Appendix, we can express $g_{\phi^{*}}$ with $g_{\rm{ref}}$ and thereby derive KL divergence. Please refer to Appendix C.1 for detailed proof. ∎
Our provided Theorem 2 guarantees the closer sampling to $p_{\rm{data}}$ or ideal MDM $\theta^{*}$ than $g_{\rm{ref}}$. Consequently, under the assumptions of two theorems, initializing with $g_{\phi_{0}}=g_{\rm ref}$ and repeatedly optimizing $g_{\phi}$ using Eq. (12) yields a convergent policy $g_{\phi^{\star}}$ that (i) attains a higher expected reward than $g_{\rm ref}$ and (ii) generates samples that are closer to the real data distribution than those from $g_{\rm ref}$.
3.4 Practical realization of unmasking policy optimization
Here, we provide practical and tractable training objectives with various $g_{\rm{ref}}$. In a real-world setting where $L$ and $|\mathcal{V}|$ are high, Eq. (12) is not directly learnable since $p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})$ is intractable in MDMs. It requires marginalizing over every trajectory to ${\bm{x}}_{0}$, i.e., $p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})=\sum_{{\bm{x}}_{1},...,{\bm{x}}_{L}}p_{\phi}({\bm{x}}_{0:L}|{\mathbf{q}})$, making both the main objective (reward maximization) and the KL term in Eq. (12) intractable. To address this, we provide a surrogate loss that can alternate the main objective.
Proposition 1 (Output–Token Level Gradient Alignment (informal)).
For MDM $\pi_{\theta}$ and unmasking policy model $g_{\phi}$ where $p_{g_{\phi}}({\bm{x}}_{n-1}\mid{\bm{x}}_{n})=g_{\phi}(a_{n}\mid{\bm{x}}_{n})\cdot\pi_{\theta}({\bm{x}}_{n-1}\mid{\bm{x}}_{n},a_{n})$ and $p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})=\sum_{{\bm{x}}_{1},...,{\bm{x}}_{L}}p_{\phi}({\bm{x}}_{0:L}|{\mathbf{q}})$, consider the output level loss $\mathcal{L}_{\rm{output}}$ and token-wise policy loss $\mathcal{L}_{\rm{token}}$:
$\mathcal{L}_{\rm{output}}(\phi)$ $=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{{\bm{x}}_{0}\sim p_{g_{\phi_{\rm{old}}}}(\cdot\mid{\mathbf{q}})}\bigr[\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0}\mid{\mathbf{q}})}A({\mathbf{q}},{\bm{x}}_{0})\bigl],$ (18)
$\mathcal{L}_{\rm{token}}(\phi)$ $=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{\{a_{1:L},{\bm{x}}_{0:L}\}\sim p_{g_{\phi_{\rm{old}}}}(\cdot\mid{\mathbf{q}})}\bigr[\sum_{n=1}^{L}\frac{g_{\phi}(a_{n}\mid{\bm{x}}_{n})}{g_{\phi_{\rm old}}(a_{n}\mid{\bm{x}}_{n})}A({\mathbf{q}},{\bm{x}}_{0})\bigl].$ (19)
Assume that we optimize $\phi$ carefully such that $\phi\approx\phi_{old}$, the two gradients are approximately equal:
$\nabla_{\phi}\mathcal{L}_{\rm{token}}\approx\nabla_{\phi}\mathcal{L}_{\rm{output}},$ (20)
and $\nabla_{\phi}\mathcal{L}_{\rm{token}}=\nabla_{\phi}\mathcal{L}_{\rm{output}}$ at first optimization step where $\phi=\phi_{\rm{old}}$.
Proof sketch..
Although the result can be readily inferred from standard RL, we provide a rigorous proof for our setting: sparse-reward, finite-horizon, episodic, and non-stationary MDPs where we maximize a GRPO-style objective. Please refer to Appendix C.2 for the full proof. ∎
Figure 2: The structure of our unmasking policy model. The model is composed of a single Transformer layer and a 3-layer MLP. In inference time, the model proceeds as follows: 1) Given an input sentence, we run the base MDMs transformer and extract a feature. 2) This feature branches into two paths: first, the base MDMs continue to produce a prediction of token distribution, and second, fed to the policy model’s transformer layer. 3) We concatenate the extracted feature with the base MDM’s Top-K probabilities. 4) The concatenated feature is then processed by a 3-layer MLP to yield a policy over unmasking positions. The shared structure with frozen MDM supports memory-efficient training (e.g., LLaDA: 8B, policy model: 134M), where we only update the unmasking policy model. The details and memory-efficient training algorithm are provided in Appendix B.2
While Eq. (18) optimizes $p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})=\sum_{{\bm{x}}_{1},...,{\bm{x}}_{L}}p_{\phi}({\bm{x}}_{0:L}|{\mathbf{q}})$ which is intractable, our surrogate loss in Eq. (19) optimizes $g_{\phi}(a_{n}\mid{\bm{x}}_{n})/g_{\phi_{\rm old}}(a_{n}\mid{\bm{x}}_{n})$ which is tractable once trajectory is sampled. Now, we detail how we train the policy model $\phi$ efficiently. Our unmasking policy model structure is shown in Figure 2, and the full training algorithm table can be found in Appendix B.2.
Final Training Objective. By using Proposition 1, clipping, and Monte Carlo approximation, we transform the theoretical loss (Eq. (12)) into tractable unmasking policy optimization (UPO) loss as follows:
$\mathcal{L}_{\rm{UPO}}(g_{\phi},g_{\rm{ref}},D)=\mathbb{E}_{{{\mathbf{q}}\sim\rho_{Q}},\{{\bm{x}}_{0:L}^{(g)},a_{1:L}^{(g)}\}_{g=1}^{G}\sim p_{g_{\phi_{\rm{old}}}}}\Biggl[\frac{1}{G}\sum_{g}^{G}\Biggl(\Bigl(\frac{1}{L}\sum_{n=1}^{L}\min\bigl(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}A_{g},$
$\text{clip}(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})},1-\epsilon,1+\epsilon)A_{g}\bigr)\Bigr)-\beta\cdot D(p_{g_{\phi}}||p_{g_{\rm{ref}}})\Biggr)\Biggr],$ (21)
where $D(p_{g_{\phi}}||p_{g_{\rm{ref}}})$ should be empirical and tractable estimator of $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$ and clipped loss has been proven to attain global optimality (Huang et al., 2024). Please note that $(g)$ in superscript of Eq. (21) is the group index.
Table 1: Various realization of $\mathcal{L}_{\rm{UPO}}(g_{\phi},g_{\rm{ref}},D)$ varying $g_{\rm{ref}}$.
Reference policy $g_{\rm{ref}}$ Max-Confidence $g_{\rm conf}$ Softmax Realization $g_{\rm conf}^{\tau}$ Top-K Realization $g_{\rm Top\text{-}K}$ Divergence $D(p_{g_{\phi}}\!||p_{g_{\rm ref}})$ Cross-Entropy Loss $\mathcal{L}_{\rm KL}({\bm{x}}_{0:L},a_{1:L},{\mathbf{q}})={\rm StopGrad}\!\Bigl(\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0:L}|{\mathbf{q}})}\cdot\bigl(1+\log\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\rm ref}}({\bm{x}}_{0:L}|{\mathbf{q}})}\bigr)\Bigr)\cdot\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ Parametrization $g_{\phi}(a^{i}\mid{\bm{x}})$ $\frac{\exp({[h_{\phi}({\bm{x}})]_{a^{i}}})}{\sum_{j\in\{\text{All Mask indices}\}}\exp({[h_{\phi}({\bm{x}})]_{a^{j}}})}$ $\frac{\exp({[h_{\phi}({\bm{x}})]_{a^{i}}})}{\sum_{j\in\{\text{Top-K Mask indices}\}}\exp({[h_{\phi}({\bm{x}})]_{a^{j}}})}$ Model initialization $\phi_{0}$ Model Pretrained with Cross-Entropy Loss w.r.t. $g_{\rm{conf}}$ Random Initialization
[…]
4 Experimental results
4.1 Experimental Setting
Models and benchmarks. We evaluate on the widely used large-scale MDM LLaDA-8B-Instruct (Nie et al., 2025) and four datasets: (i) logic-puzzle benchmarks—Sudoku (Zhao et al., 2025) and Zebra (Shah et al., 2024)—and (ii) mathematical-reasoning benchmarks—GSM8K (Cobbe et al., 2021) and Math500 (Lightman et al., 2023).
Generation protocols and rewards. We utilize LLaDA-8B-Instruct (Nie et al., 2025) for all experiments. For logic puzzles, the model receives a prompt and a masked solution; it then denoises by unmasking one position at a time. The reward is the fraction of slots predicted correctly (e.g., in Sudoku, the proportion of blank cells filled with the correct digits). For mathematical reasoning, the model is given a problem and denoises a length-128 masked sequence. Following the LLaDA generation protocol, we partition the 128 positions into four bins of size 32 and perform any-order sampling sequentially over the bins. The reward is $1$ if the final answer matches the ground truth and 0 otherwise.
4.2 Training dynamics across reference policies
Figure 3: Training accuracy on Sudoku and GSM8K under $\mathcal{L}_{\rm UPO}(g_{\phi},g_{\rm ref},D)$ for different choices of $g_{\rm ref}$. Reference realizations are summarized in Table 1. “Random initialization” denotes training a randomly initialized model with $\mathcal{L}_{\rm UPO}(g_{\phi},\varnothing,\varnothing)$.
We train our policy model on GSM8K and Sudoku by optimizing $\mathcal{L}_{\rm UPO}(g_{\phi},g_{\rm ref},D)$ with three reference realizations from Table 1: $g_{\rm conf}$, $g_{\rm conf}^{\tau}$, and $g_{\rm Top\text{-}K}$. We also report ablation that drops the divergence term, i.e. $\mathcal{L}_{\rm UPO}(g_{\phi},\varnothing,\varnothing)$, with a randomly initialized policy, which is conceptually equivalent to DColT. The result is shown in Figure 3. The strongest configuration depends on the task: on Sudoku, max-confidence realization achieves the best accuracy, while on GSM8K, Top-$K$ realization performs best. We conjecture that in GSM8K, restricting exploration to the Top–$K$ indices stabilizes learning under $\mathcal{L}_{\rm UPO}$ due to the much larger number of masked tokens. Across the two benchmarks, our optimal configuration outperforms the randomly initialized model—that is, DColT (green line)—by margins of at least 2% and 3%, respectively.
4.3 Main result
Table 2: Accuracy of each sampling strategy. Best per row in bold.
Benchmark Sampler Random (Zheng et al., 2025) Margin (Kim et al., 2025) Entropy (Ye et al., 2025) Confidence (Chang et al., 2022) Ours SUDOKU $0.616$ $0.713$ $0.671$ $0.705$ 0.817 Zebra $0.339$ $0.346$ $0.351$ $0.337$ 0.362 GSM8K $0.612$ $0.671$ $0.667$ $0.684$ 0.703 Math500 $0.196$ 0.284 $0.266$ $0.272$ 0.284
Baselines and setup. We compare our learned unmasking policy with four widely used schedulers: random (Zheng et al., 2025), margin (Kim et al., 2025), entropy (Ye et al., 2025), and confidence (Chang et al., 2022). For Sudoku and Zebra, we utilize max-confidence realization, and for GSM8K and Math500, we use Top-$K$ realization where $K=5$.
Results. Table 2 shows that our policy matches or surpasses all heuristics on every benchmark. For Sudoku and Zebra, where we use $g_{\rm conf}$ as the reference, our model exceeds this reference: Sudoku improves from 70.5% to 81.7%, and Zebra from 33.7% to 36.2%. On GSM8K, trained with a Top-$K$ reference, our policy reaches 70.3% and exceeds the strong max-confidence baseline at 68.4%. On Math500, also trained with a Top-$K$ reference, our policy attains 28.4%, tying the best baseline while improving over max-confidence at 27.2%. Overall, max-confidence remains a strong single-path heuristic, yet a learned policy trained in our regularized framework consistently surpasses it.
4.4 Ablation Study
Table 3: Accuracy on GSM8K of post-trained LLaDA with diffu-GRPO by sampling strategy.
Random (Zheng et al., 2025) Margin (Kim et al., 2025) Entropy (Ye et al., 2025) Confidence (Chang et al., 2022) Ours $0.638$ $0.750$ $0.740$ $0.751$ 0.764
Figure 4: Mean and standard deviation of group reward during training under $\mathcal{L}_{\rm UPO}$. For SUDOKU, we compare pretrained policy trained with/without CE to $g_{\rm conf}$, and for GSM8K, we compare randomly initialized $g_{\phi_{\rm Top\text{-}K}}$ trained with/without KL to $g_{\rm Top\text{-}K}$.
Compatibility with diffu-GRPO. Zhao et al. (2025) introduces diffu-GRPO, an RL method that post-trains the MDM itself via GRPO. Our approach also uses RL but targets a different component: we optimize the unmasking policy while keeping the base MDM fixed. The two methods are therefore complementary and can be combined. To demonstrate this, we first post-train LLaDA-8B-Instruct on GSM8K with diffu-GRPO, then train our policy model with Top-$K$ realization, $\mathcal{L}_{\rm{UPO}}(g_{\phi_{\rm{Top-5}}},g_{\rm{Top-5}},\mathcal{L}_{\rm{KL}})$. As reported in Table 3, our unmasking policy model yields additional gains on top of the diffu-GRPO model: +12.6% over random and +1.3% over max-confidence.
Power of regularization. We compare the mean and standard deviation of the group reward while optimizing $\mathcal{L}_{\rm UPO}(g_{\phi},g_{\rm ref},D)$ in two setups: for Sudoku, we train a pretrained policy with/without CE to $g_{\rm conf}$, and for GSM8K, we train a randomly initialized $g_{\phi_{\rm Top\text{-}K}}$ with/without $\mathcal{L}_{\rm KL}$ to $g_{\rm Top\text{-}K}$. As shown in Figure 4, adding the divergence term yields higher final accuracy and, at similar mean rewards—even near 1—maintains a larger group-reward standard deviation. This helps the policy model avoid premature convergence under the GRPO framework and reach a higher convergence point. We conjecture that, for GSM8K, $\mathcal{L}_{\rm KL}$ to $g_{\rm Top\text{-}K}$ force the probability of the trajectory sampled from $g_{\phi_{\rm{Top-K}}}$ to be all equal and prevent early path collapse. For Sudoku, we suppose that CE toward $g_{\rm conf}$ counteracts collapse onto a suboptimal order by pulling the policy back to high-confidence alternatives, keeping multiple competitive paths active and maintaining variance.
5 Conclusion
We introduced a KL-regularized MDP formulation for learning the unmasking policy in masked diffusion models. Our analysis provides two guarantees: convergence to a fixed point that improves expected reward over the reference policy, and a KL tightening showing that the terminal-output distribution is closer to the real data than the reference. To make the framework practical, we derived a tractable surrogate loss and proposed several reference realizations. Across Sudoku, Zebra, GSM8K, and Math500, the learned unmasking policy model consistently matches or surpasses strong heuristics.
Future directions. On Sudoku, max-confidence exceeds Top-$K$, suggesting that near-optimal training can benefit from occasional low-confidence selections. Meanwhile, on GSM8K, the much larger search space makes Top-$K$ preferable for stability. In future work, we plan to develop an unmasking policy optimization algorithm that explores the full mask tokens even when the search space is large.
Appendix A Related works
Discrete Diffusion Models. Diffusion probabilistic models have become the state-of-the-art approach for continuous data domains such as images, audio, and video, by iteratively denoising Gaussian-corrupted inputs (Ho et al., 2020; Song et al., 2021). Extending this success to discrete data (e.g. text or sequences) has led to the development of discrete diffusion models built on discrete-state Markov chains rather than continuous noise processes (Austin et al., 2021; Hoogeboom et al., 2021). Lou et al. (2024) proposed SEDD, which utilizes theoretical score entropy objective. Meanwhile, masked diffusion models (MDMs), which utilize an absorbing mask strategy, have emerged as a leading class of discrete diffusion models for text generation (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024). By using an absorbing masking state, MDMs simplify the corruption process: once a token is masked, it remains masked until generation, analogous to absorbing diffusion in continuous models (Shi et al., 2024). This simplification yields a clean and principled training recipe aligned with continuous-time diffusion theory, including an evidence lower bound (ELBO) objective (Shi et al., 2024). Beyond continuous-time discrete diffusion, Zheng et al. (2025); Ou et al. (2025) provides a time-agnostic training framework and simple cross-entropy loss. These advances have significantly closed the performance gap between MDMs and traditional autoregressive models on language tasks (Ye et al., 2025; Nie et al., 2025) when the scale of the model is similar.
Unmasking Strategy in Masked Diffusion Inference. A crucial factor for MDM performance is the unmasking strategy used during inference—i.e., how the model decides which masked tokens to reveal at each generation step. While MDMs have shown superior performance on text generations, earlier works (Ou et al., 2025; Shih et al., 2022; Li et al., 2021) have empirically observed that a random unmasking strategy is suboptimal. Moreover, Kim et al. (2025) proves that no polynomial-time algorithm can solve any-order generation; i.e., we cannot train MDMs that exactly recovers the real data distribution for all masked sentences. However, they empirically show that selecting unmasking tokens via max-confidence (Chang et al., 2022) or max-margin (Kim et al., 2025) at inference time can bypass sub-hard instances and even outperform ARMs. As a result, contemporary large-scale MDMs (Nie et al., 2025; Ye et al., 2025) commonly adopt max-confidence scheduling at inference.
Figure 5: The structure of our unmasking policy model. The model is composed of a single Transformer layer and a 3-layer MLP. In inference time, the model proceeds as follows: 1) Given an input sentence, we run the base MDMs transformer and extract a feature. 2) This feature branches into two paths: first, the base MDMs continue to produce a prediction of token distribution, and second, fed to the policy model’s transformer layer. 3) We concatenate the extracted feature with the base MDM’s Top-K probabilities. 4) The concatenated feature is then processed by a 3-layer MLP to yield a policy over unmasking positions.
Appendix B Training recipe in detail
In this section, we explain our memory-efficient training recipe in detail. We provide full derivations of various reference policy realizations (Appendix B.1), and algorithms for training and generating (Appendix B.2). For a deeper understanding of our memory-efficient training algorithm and model architecture, we first provide notations and model inference in detail.
Policy model inference process.
Our rationale behind policy model structure is two: 1) utilizing the MDM’s knowledge (extracted feature), which allows for avoiding from-scratch training and allow memory-efficient training, and 2) utilizing the MDM’s token prediction distribution, which can help the model mimic the behavior of $g_{\rm{conf}}$. The main paper outlined the policy model’s inference at a high level; here, we provide a notation and description of how, given a masked sentence ${\bm{x}}_{n}$, the model generates an unmasking trajectory. For reference, we reproduce the model architecture from the main paper (Figure 2) as Figure 5. To recap, the unmasking policy model is parameterized as a softmax over masked positions using a learnable scorer $h_{\phi}({\bm{x}})\in\mathbb{R}^{L}$:
$g_{\phi}(a^{i}\,|\,{\bm{x}}_{n})=\frac{\exp\!\big([h_{\phi}({\bm{x}}_{n})]_{a^{i}}\big)}{\sum_{i^{\prime}=1}^{n}\exp\!\big([h_{\phi}({\bm{x}}_{n})]_{a^{i^{\prime}}}\big)},$ (22)
where $a^{i}$ is $i$-th mask of ${\bm{x}}_{n}$. As illustrated in Figure 5, the policy model decomposes as $\phi=(\phi_{\rm MLP},\phi_{\rm TF})$, where $\phi_{\rm TF}$ is a 1-layer Transformer and $\phi_{\rm MLP}$ is a 3-layer MLP. We similarly write the MDM as $\theta=(\theta_{\rm TF},\theta_{\rm MLP})$ with a feature extractor transformer $\theta_{\rm TF}$ and a token prediction head $\theta_{\rm MLP}$. Given the current masked sequence ${\bm{x}}_{n}$ and selected mask $a^{i}$, the MDM computes the token distribution as follows:
$\mathbf{H}_{n}=\theta_{\rm TF}({\bm{x}}_{n})\in\mathbb{R}^{L\times d},\qquad\mathbf{u}_{a^{i}}=\mathbf{H}_{n}[a^{i}]\in\mathbb{R}^{d},$
$\mathbf{z}_{a^{i}}=\theta_{\rm MLP}(\mathbf{u}_{a^{i}})\in\mathbb{R}^{|\mathcal{V}|},\qquad\pi_{\theta}\!\left(x^{a^{i}}_{n-1}=c\,\middle|\,{\bm{x}}_{n}\right)=\mathrm{softmax}(\mathbf{z}_{a^{i}})_{c}\;\;\;\;(c\in\mathcal{V}),$
where $d$ is the output dimension of the transformer feature. Let $\mathbf{p}_{a^{i}}=\mathrm{softmax}(\mathbf{z}_{a^{i}})$ the token posterior at index $a^{i}$ with dimension $|\mathcal{V}|$. The policy encoder first refines the sequence features,
$\widetilde{\mathbf{H}}_{n}=\phi_{\rm TF}(\mathbf{H}_{n})\in\mathbb{R}^{L\times d},\qquad\widetilde{\mathbf{u}}_{a^{i}}=\widetilde{\mathbf{H}}_{n}[a^{i}]\in\mathbb{R}^{d}.$
Let $\mathrm{TopK}(\mathbf{p}_{a^{i}})\in\mathbb{R}^{K}$ denote the vector of the $K$ largest probabilities of $\mathbf{p}_{a^{i}}$, sorted in descending order. We form the concatenated feature
$\mathbf{s}_{a^{i}}=\big[\,\widetilde{\mathbf{u}}_{a^{i}}\;;\;\mathrm{TopK}(\mathbf{p}_{a^{i}})\,\big]\in\mathbb{R}^{d+K},$ (23)
and map it to a scalar score via the 3-layer MLP:
$h_{\phi}({\bm{x}}_{n})]_{a^{i}}=\phi_{\rm MLP}(\mathbf{s}_{a^{i}})\in\mathbb{R}.$
Finally, the unmasking policy over the masked index set is the softmax of these scores (Eq. (22)).
B.1 Full Derivations for $\mathcal{L}_{\rm UPO}$ Realizations
General formulations.
To recap, our final objective is given as follows:
$\mathcal{L}_{\rm{UPO}}(g_{\phi},g_{\rm{ref}},D)=\mathbb{E}_{{{\mathbf{q}}\sim\rho_{Q}},\{{\bm{x}}_{0:L}^{(g)},a_{1:L}^{(g)}\}_{g=1}^{G}\sim p_{g_{\phi_{\rm{old}}}}}\Biggl[\frac{1}{G}\sum_{g}^{G}\Biggl(\Bigl(\frac{1}{L}\sum_{n=1}^{L}\min\bigl(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}A_{g},$
$\text{clip}(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})},1-\epsilon,1+\epsilon)A_{g}\bigr)\Bigr)-\beta\cdot D(p_{g_{\phi}}||p_{g_{\rm{ref}}})\Biggr)\Biggr],$ (24)
The divergence term $D$ in Eq. (24) is a functional that, given a prompt $\mathbf{q}$}, a full trajectory $\tau=({\bm{x}}_{0:L},a_{1:L})$, and models $(g_{\phi},g_{\rm ref},\pi_{\theta})$, returns a scalar estimate. Formally,
$D_{g_{\phi},g_{\rm ref},\pi_{\theta}}:\;\mathcal{Q}\times\mathcal{X}_{0:L}\times\mathcal{A}_{1:L}\rightarrow\mathbb{R},$
and its empirical expectation serves as a tractable estimator of the output-level KL divergence $D_{\rm KL}\!\big(p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})\,\|\,p_{g_{\rm ref}}({\bm{x}}_{0}\mid{\mathbf{q}})\big)$. Following Proposition 2 and Proposition 3, we can decrease the output-level KL divergence with $\mathcal{L}_{\rm{KL}}$ defined as follows:
$\mathcal{L}_{\rm KL}({\bm{x}}_{0:L},a_{1:L},{\mathbf{q}})={\rm StopGrad}\!\Bigl(\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0:L}|{\mathbf{q}})}\cdot\bigl(1+\log\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\rm ref}}({\bm{x}}_{0:L}|{\mathbf{q}})}\bigr)\Bigr)\cdot\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n}),$ (25)
where $g_{\phi},g_{\rm ref}$ and $\pi_{\theta}$ is omitted in $\mathcal{L}_{\rm{KL}}$ for brevity. We can convert $p_{g}({\bm{x}}_{0:L}|{\mathbf{q}})$ into $\prod_{n=1}^{L}g(a_{n}|{\bm{x}}_{n})\cdot\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n},{\mathbf{q}})$ in Eq. (25):
$\mathcal{L}_{\rm KL}(\cdot)={\rm StopGrad}\!\Bigl(\prod_{n=1}^{L}\frac{g_{\phi}(a_{n}|{\bm{x}}_{n})}{g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})}\cdot\bigl(1+\sum\log\frac{g_{\phi}(a_{n}|{\bm{x}}_{n})}{g_{{\rm{ref}}}(a_{n}|{\bm{x}}_{n})}\bigr)\Bigr)\cdot\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n}).$ (26)
We already computed $g_{\phi_{\rm{old}}}(\cdot)$ while sampling, and $g_{\phi}(\cdot)$ while computing reward-maximization objective term in $\mathcal{L}_{\rm{UPO}}$, and $g_{\rm{ref}}$ does not require network evaluations. Therefore, we can measure $\mathcal{L}_{\rm{KL}}$ without no additional computation. Furthermore, since $\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ is summation, we can perform gradient update token-by-token, which allows tractable and memory-efficient training.
Softmax realization.
For softmax realization of $g_{\rm{ref}}$, we consider $g_{\rm{conf}^{\tau}}$ given as follows:
$g_{\rm{conf}^{\tau}}(a^{i}|{\bm{x}}_{n})$ $=\frac{\sum_{c\in\mathcal{V}}\exp({\pi_{\theta}(x_{n-1}^{a^{i}}=c|{\bm{x}}_{n})/\tau})}{\sum_{a^{j}}\sum_{c\in\mathcal{V}}\exp({\pi_{\theta}(x_{n-1}^{a^{j}}=c|{\bm{x}}_{n})/\tau})},\;\;\text{s.t.}\;\lim_{\tau\rightarrow 0+}g_{\rm{conf}^{\tau}}=g_{\text{conf}}.$ (27)
Here, since $g_{\rm{conf}^{\tau}}$ assigns non-zero probability to every unmasking index, we can estimate the terminal KL divergence with a sampled trajectory. We utilize $\mathcal{L}_{\rm{KL}}$ (Eq. (26)) by replacing $g_{\rm{ref}}$ with $g_{\rm{conf}^{\tau}}$.
Top-$K$ realization
For Top-$K$ realization of $g_{\rm{ref}}$, we consider $g_{\rm{conf}^{\tau}}$ given as follows:
$g_{\rm{Top-K}}(a^{i}|{\bm{x}}_{n})$ $=\frac{1}{K}\cdot\textbf{1}(a^{i}\in\underset{{a^{j}}}{\rm{argtopk}}(\max_{c\in\mathcal{V}}\pi_{\theta}({\bm{x}}_{n-1}^{a^{j}}=c|{\bm{x}}_{n}))),\;\;\text{s.t.}\;g_{\rm{Top-1}}=g_{\rm{conf}}.$ (28)
where we reparametrize $g_{\phi_{\rm{Top-K}}}$ defined as follows:
$g_{\phi_{\rm{Top-K}}}(a^{i}|{\bm{x}}_{n})=\frac{\exp({[h_{\phi}({\bm{x}}_{n})]_{a^{i}}})}{\sum_{a^{j}\in{\rm{argtopk}}\max\pi_{\theta}(\cdot)}\exp({[h_{\phi}({\bm{x}}_{n})]_{a^{j}}})},$ (29)
which is the softmax distribution of $h_{\phi}$ on the Top-$K$ set of unmasking positions. This reparametrization offers two advantages: 1) when $\phi$ is randomly initialized, we can expect $g_{\phi_{\rm{Top-K}}}=g_{\rm{Top-K}}$, and 2) for sampled trajectory, $g_{\rm{Top-K}}(a_{n}|{\bm{x}}_{n})$ is non-zero for all $n\in[1,L]$ since we sample from $g_{\phi_{\rm{Top-K}}}$. Therefore, we can train $g_{\phi_{\rm{Top-K}}}$ from scratch with $\mathcal{L}_{\rm{KL}}$ by replacing $g_{\rm{ref}}=g_{\rm{Top-K}}$.
Max-confidence realization.
For max-confidence realization of $g_{\rm{ref}}$, we consider $g_{\rm{conf}}$ given as follows:
$g_{\text{conf}}(a^{i}|{\bm{x}}_{n})=\textbf{1}\bigl(a^{i}=\operatorname*{arg\,max}_{a^{j}}(\max_{c\in\mathcal{V}}\pi_{\theta}({\bm{x}}_{n-1}^{a^{j}}=c|{\bm{x}}_{n}))\bigr),$ (30)
where $g_{\rm{conf}}$ is deterministic. The terminal KL divergence $D_{\rm KL}\!\big(p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})\,\|\,p_{g_{\rm ref}}({\bm{x}}_{0}\mid{\mathbf{q}})\big)$ can be defined where $\pi_{\theta}$ is stochastic, but we cannot empirically estimate the terminal KL divergence with sampled trajectory since there exists trajectory sampled from $p_{g_{\phi}}$ is 0 for $p_{g_{\rm{conf}}}$. Therefore, we just utilize cross-entropy given as follows:
${\rm{CE}}_{g_{\phi},g_{\rm{conf}},\pi_{\theta}}({\bm{x}}_{0:L},{\mathbf{q}})=\sum_{n=1}^{L}\sum_{a\in\mathcal{A}_{{\bm{x}}_{n}}}g_{\rm{conf}}(a|{\bm{x}}_{n})\log g_{\theta}(a|{\bm{x}}_{n}).$ (31)
Although not theoretical, we can expect to regularize $g_{\phi}$ being far away from $g_{\rm{conf}}$, and we have empirically confirmed the performance gain (Section 4.4).
B.2 Algorithms
Algorithm 1 Generalized Sampling of MDMs
1:Sequence length $L$, vocabulary $\mathcal{V}=\{1,\dots,m\}$, pretrained masked diffusion model $\pi_{\theta}$, unmasking policy $g$
[…]
3:for $n\leftarrow L\ \textbf{to}\ 1$ do
4:  sample $a_{n}\sim g(a|{\bm{x}}_{n})$ $\triangleright$ unmasking token distribution. e.g., $\mathcal{U}(\{a^{i}\}_{i=1}^{n})$
5:  sample $x_{n-1}^{a_{n}}\sim\pi_{\theta}(x_{n-1}^{a_{n}}|{\bm{x}}_{n})$ $\triangleright$ per-position categorical probabilities
6:  ${\bm{x}}_{n-1}\leftarrow\{x_{n}^{1},x_{n}^{2},\dots,x_{n}^{a_{n}-1},x_{n-1}^{a_{n}},x_{n}^{a_{n}+1},\dots,x^{L}_{n}\}$
7:end for
8:Output: ${\bm{x}}_{0}$
Algorithm 2 Memory-efficient training of the unmasking policy model
1:Pretrained masked diffusion model $\pi_{\theta}$ composed of feature extracting transformer $\theta_{\rm{TF}}$ and token estimator $\theta_{\rm{MLP}}$, unmasking policy model $g_{\phi}$ composed of 1-layer transformer $\phi_{\rm{TF}}$ and 3-Layer MLP $\phi_{\rm{MLP}}$, divergence loss $D$, reference policy $g_{\rm{ref}}$, prompt distribution $\mathcal{Q}$, number of inner updates $\mu$
2:while not converge do
3:Sampling phase
4:  with torch.no_grad():
5:   $g_{\phi_{\rm{old}}}\leftarrow g_{\phi}$
6:   Sample a prompt ${\mathbf{q}}\sim\mathcal{Q}$
7:   Sample G trajectories $\{{\bm{x}}_{0:L}^{(g)},a_{1:L}^{(g)}\}_{g=1}^{G}\sim p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})$, $g\in[G]$
8:   $\triangleright$ while sampling, save ${\mathbf{H}}_{n}^{(g)}=\theta_{\rm{TF}}({\bm{x}}_{n}^{(g)})$, $\mathcal{A}_{{\bm{x}}_{n}^{(g)}}$, $\pi_{\theta}(\cdot|{\bm{x}}_{n}^{(g)})$, $g_{\phi_{old}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})$
9:   $\triangleright$ if $D=\mathcal{L}_{\rm{KL}}$, record $p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}^{(g)},a_{1:L}^{(g)}|{\mathbf{q}})$, $p_{g_{\phi_{\rm{ref}}}}({\bm{x}}_{0:L}^{(g)},a_{1:L}^{(g)}|{\mathbf{q}})$ while sampling
10:   For each ${\bm{x}}_{0}^{(g)}$, compute reward $r_{g}$ and advantage $A_{g}$
11:  end with
12:Gradient update phase
13:  for gradient update iterations $n=1,\dots,\mu$ do
14:   with torch.no_grad():
15:     For each ${\bm{x}}_{0:L}^{(g)}$, pre-compute $p_{g_{\phi}}({\bm{x}}_{0:L}^{(g)}|{\mathbf{q}})$
16:     if $D=\mathcal{L}_{\rm{KL}}$ then
17:      $\kappa_{g}\leftarrow\Bigl(\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0:L})}\cdot\bigl(1+\log\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L})}{p_{g_{\rm ref}}({\bm{x}}_{0:L})}\bigr)\Bigr)$
18:     end if
19:   end with
20:   Make mini-batch $\{B\}_{i=1}^{b}$ such that $B_{1}\cup B_{2}\cup\dots\cup B_{b}=\{1,L\}$ and $B_{i}\cap B_{j}=\emptyset$ for all $i\neq j$
21:   for $B\in\{B_{i}\}_{i=1}^{b}$ do
22:     For each $n\in B$ and $g\in\{1,..,G\}$, $\widetilde{{\mathbf{H}}}_{n}^{(g)}\leftarrow\phi_{\rm{TF}}({\mathbf{H}}_{n}^{(g)})$
23:     For each $n\in B$ and $g\in\{1,..,G\}$, construct ${\mathbf{s}}_{n}^{(g)}$ using Eq. (23)
24:     $g_{\phi}(\cdot|{\bm{x}}_{n}^{(g)})\leftarrow\phi_{\rm{MLP}}({\mathbf{s}}_{n}^{(g)})$
25:     $\mathcal{L}_{\phi}\leftarrow\frac{1}{|B|}\sum_{n\in B,g\in G}\min\bigl(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}A_{g},\text{clip}(\frac{g_{\phi}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})}{g_{\phi_{\rm{old}}}(a_{n}^{(g)}|{\bm{x}}_{n}^{(g)})},1-\epsilon,1+\epsilon)A_{g}\bigr)$
26:     if $D=\rm{CE}$ then
27:      $\mathcal{L}_{\phi}\leftarrow\mathcal{L}_{\phi}-\beta\cdot\sum_{g\in G}\sum_{n\in B}{\rm{CE}}(g_{\rm{conf}}(\cdot|{\bm{x}}_{n}^{(g)})||g_{\phi}(\cdot|{\bm{x}}_{n}^{(g)}))$
28:     end if
29:     if $D=\mathcal{L}_{\rm{KL}}$ then
30:      $\mathcal{L}_{\phi}\leftarrow\mathcal{L}_{\phi}-\beta\cdot\sum_{g\in G}\kappa_{g}\sum_{n\in B}\log g_{\phi}(a_{n}^{(g)}\mid{\bm{x}}_{n}^{(g)})$
31:     end if
32:     Update $\phi$ by performing gradient ascent on $\mathcal{L}_{\rm{UPO}}$
33:   end for
34:  end for
35:end while
36:Output: $g_{\phi}$
Sampling algorithm.
We provide generalized sampling algorithm of large-scale MDMs in Algorithm 1. The sampling trajectory differs by unmasking policy $g$, and we can perform various sampling by setting $g$ as $g_{\rm{rand}}$, $g_{\rm{conf}}$, $g_{\rm{conf}^{\tau}}$, $g_{\rm{Top-K}}$, or $g_{\phi}$.
Unmasking policy training algorithm.
We provide a memory-efficient unmasking policy model training algorithm in Algorithm 2. We have previously mentioned how we utilize the extracted features and token distribution of MDM as input to the unmasking policy model. We also have provided how the gradient of $\mathcal{L}_{\rm{UPO}}$ can be measured by the token-wise gradient of $g_{\phi}(\cdot|{\bm{x}})$. Summing up all techniques, we can perform training without using GPU memory more than the amount of original MDM occupies. Indeed, we have confirmed that training can be performed with only one A100 40GB GPU.
Specifically, the algorithm is divided into two phases: the sampling phase and the gradient update phase. At the sampling phase, we sample various sequences from $p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})$. In this phase, we gather the extracted feature of the MDM transformer, the token distribution of MDM, the mask token index, and record the probability of the trajectory. At the gradient update phase, we can perform mini-batch training where all the gradients of $g_{\phi}(\cdot|{\bm{x}}_{n})$ in $\mathcal{L}_{\rm{UPO}}$ are independent of others. Furthermore, since we have gathered all the information of $\pi_{\theta}$ that is needed to update the gradient in the sampling phase, we can free the memory of MDM $\theta$. The only parameters requiring forward and backward process in the gradient update phase is the unmasking policy model, which is much smaller than MDM (e.g., LLaDA: 8B, corresponding policy model: 134M). Therefore, we can perform memory-efficient training. For more details, please refer to the algorithm table.
Appendix C Detailed Proof
In this section, we provide full proofs omitted in the main paper.
C.1 proof of KL tightening for MDM policy
See 2
Proof.
We first recall the GRPO policy training dynamics learning Eq. (12) proven by Mroueh (2025):
Lemma 1 (GRPO Policy Dynamic (Mroueh, 2025)).
Optimal GRPO iterations policies solving Eq. (12) satisfy the following recursion, for $n\geq 1$:
$p_{g_{\phi_{n}}}({\bm{x}}_{0}|{\mathbf{q}})$ $=\frac{1}{Z_{n-1}({\mathbf{q}})}p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})$
$\cdot\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))\cdot r({\mathbf{q}},{\bm{x}}_{0}))-w_{\epsilon}^{-}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))\cdot(1-r({\mathbf{q}},{\bm{x}}_{0}))))$ (32)
where
$Z_{n-1}({\mathbf{q}})=r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(-\frac{1}{\beta}(w_{\epsilon}^{-}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))$ (33)
and $w_{\epsilon}^{+}(r)$ and $w_{\epsilon}^{-}(r)$ are defined as follows:
$w_{\epsilon}^{+}(r)=\frac{1-r}{r(1-r)+\epsilon},\;\;\;\;w_{\epsilon}^{-}(r)=\frac{r}{r(1-r)+\epsilon}$ (34)
Under Theorem 1, at the fixed point, we can substitute the converged policy into the recursion and compare the two KL terms. The next steps simply expand the KL difference. Then, for the convergence point where $\lim_{n\rightarrow{\infty}}r_{g_{\phi_{n}}}=r_{g^{*}}$:
$D_{\rm{KL}}$ $(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}}))-D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))$ (35)
$=\mathbb{E}_{p_{\theta^{*}}}[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})]-\mathbb{E}_{p_{\theta^{*}}}[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})]$ (36)
$=\mathbb{E}_{p_{\theta^{*}}}[\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})]$ (37)
$=\mathbb{E}_{p_{\theta^{*}}}[\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))\cdot r({\mathbf{q}},{\bm{x}}_{0}))-w_{\epsilon}^{-}(r_{g_{\phi^{*}}}({\mathbf{q}}))\cdot(1-r({\mathbf{q}},{\bm{x}}_{0}))]-\log Z_{\phi^{*}}$ (38)
$=\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))-\log Z_{\phi^{*}}\hskip 28.45274pt\because\forall{\bm{x}}_{0}\in\mathcal{X},p_{\theta^{*}}({\bm{x}}|{\mathbf{q}})>0\rightarrow r({\mathbf{q}},{\bm{x}}_{0})=1$ (39)
$=\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))$
$-\log\Biggr(r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(-\frac{1}{\beta}(w_{\epsilon}^{-}(r_{g_{\phi^{*}}}({\mathbf{q}}))\Biggl)$ (40)
$\geq\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))$
$-\log\Biggr(r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))\Biggl)$ (41)
$\hskip 170.71652pt\because\forall r\in\{0,1\},\;\;w_{e}^{+}(r)\geq 0,\;w_{e}^{-}(r)\geq 0$
$=0,$ (42)
where $\operatorname{supp}(p_{\theta^{*}}(\cdot\mid{\mathbf{q}}))\subseteq\operatorname{supp}(p_{g_{\mathrm{ref}}}(\cdot\mid{\mathbf{q}}))$. Therefore, the inequality in Eq. (16) holds. Accordingly, inequality in Eq. (17) holds where $p_{\theta^{*}}=p_{\rm{data}}$ for ideal MDM $\theta^{*}$. ∎
C.2 Surrogate loss of output-level reward maximization
Our goal here is to show that the output-level reward maximization term in Eq. (12) can be converted into a tractable token-wise reward maximization term. The conclusion of Proposition 1 aligns with conventional RL that have developed with rigorous proofs, yet in our setting, where training the MDM unmasking policy, the assumptions are slightly different. More specifically, our MDP is a finite-horizon MDP with fixed length $L$ (text length is determined), sparse reward (only a verifiable reward for the terminal state exists), and non-stationary (for different time steps, states do not overlap). To fill this gap, we rigorously show that Proposition 1 holds.
We cannot even directly apply the policy gradient theorem (Sutton et al., 1999), which is essential in the proof. Therefore, we start with justifying how the policy gradient theorem also works in our framework, which can be easily proven:
Lemma 2 (Policy gradient theorem in sparse-reward, finite-horizon, episodic, and non-stationary MDP).
Fix a horizon $L$. For each $i\in\{0,1,\dots,L\}$ let the state space be $\mathcal{X}_{n}$, and assume the spaces are pairwise disjoint: $\mathcal{X}_{n}\cap\mathcal{X}_{j}=\emptyset$ for $i\neq j$. At step $i\in\{1,\dots,L\}$, from state ${\bm{x}}_{n}\in\mathcal{X}_{n}$ a discrete action $a_{n}\in\mathcal{A}_{n}$ is drawn from a stochastic policy $g_{\phi}(a_{n}\mid{\bm{x}}_{n})$, and the environment transitions to the next layer’s state ${\bm{x}}_{n-1}\in\mathcal{X}_{n-1}$ according to dynamics $\pi({\bm{x}}_{n-1}\mid{\bm{x}}_{n},a_{n})$. Assume the start state is fixed as ${\bm{x}}_{L}$. The trajectory is $\tau=({\bm{x}}_{0:L},a_{1:L})$ with joint density
$p_{g_{\phi}}(\tau\mid{\bm{x}}_{L})=\prod_{n=1}^{L}g_{\phi}(a_{n}\mid{\bm{x}}_{n})\,\pi({\bm{x}}_{n-1}\mid{\bm{x}}_{n},a_{n}).$
Rewards are terminal-only: $r({\bm{x}}_{0})$ at $i=0$ and 0 otherwise. The objective is the expected terminal reward
$J(\phi)\;=\;\mathbb{E}_{\tau\sim p_{g_{\phi}}(\cdot\mid{\bm{x}}_{L})}\!\big[r({\bm{x}}_{0})\big].$
Define the value function and state-action value function under $g_{\phi}$ by
$V_{g_{\phi}}({\bm{x}}_{n})\;=\;\mathbb{E}\!\left[r({\bm{x}}_{0})\mid{\bm{x}}_{n}\right],\qquad Q_{g_{\phi}}({\bm{x}}_{n},a_{n})\;=\;\mathbb{E}\!\left[r({\bm{x}}_{0})\mid{\bm{x}}_{n},a_{n}\right].$
Under the above assumptions, the policy gradient depends only on the policy $g_{\phi}$:
$\nabla_{\phi}J(\phi)$ $=\;\nabla_{\phi}V_{g_{\phi}}({\bm{x}}_{L})=\sum_{{\bm{x}}\in\mathcal{X}_{1}\cup\mathcal{X}_{2}\dots\cup\mathcal{X}_{L}}p_{g_{\phi}}({\bm{x}})\sum_{a}Q_{g_{\phi}}({\bm{x}},a)\nabla_{\phi}g_{\phi}(a|{\bm{x}}).$ (43)
Proof.
By definition,
$J(\phi)=\mathbb{E}_{\tau\sim p_{g_{\phi}}(\cdot\mid{\bm{x}}_{L})}[r({\bm{x}}_{0})].$
Using the log-derivative trick,
$\nabla_{\phi}J(\phi)$ $=\sum_{\tau}r({\bm{x}}_{0})\,\nabla_{\phi}p_{g_{\phi}}(\tau\mid{\bm{x}}_{L})$
$=\sum_{\tau}r({\bm{x}}_{0})\,p_{g_{\phi}}(\tau\mid{\bm{x}}_{L})\,\nabla_{\phi}\log p_{g_{\phi}}(\tau\mid{\bm{x}}_{L}).$
Since only the policy $g_{\phi}$ depends on $\phi$, we have
$\log p_{g_{\phi}}(\tau\mid{\bm{x}}_{L})=\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})+\text{const},$
and therefore
$\nabla_{\phi}\log p_{g_{\phi}}(\tau\mid{\bm{x}}_{L})=\sum_{n=1}^{L}\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n}).$
Thus,
$\nabla_{\phi}J(\phi)$ $=\mathbb{E}_{\tau\sim p_{g_{\phi}}(\cdot\mid{\bm{x}}_{L})}\!\left[r({\bm{x}}_{0})\sum_{n=1}^{L}\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\right].$
Taking conditional expectations with respect to $({\bm{x}}_{n},a_{n})$ yields for each $i\in\{1,\dots,L\}$ we have
$\mathbb{E}_{\tau}\!\Big[r({\bm{x}}_{0})\,\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ $=\mathbb{E}_{{\bm{x}}_{n},a_{n}}\!\Big[\mathbb{E}\!\big[r({\bm{x}}_{0})\mid{\bm{x}}_{n},a_{n}\big]\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ (tower property)
$=\mathbb{E}_{{\bm{x}}_{n},a_{n}}\!\Big[Q_{g_{\phi}}({\bm{x}}_{n},a_{n})\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ (44)
$=\sum_{{\bm{x}}_{n}\in\mathcal{X}_{n}}\sum_{a_{n}\in\mathcal{A}_{n}}p_{g_{\phi}}({\bm{x}}_{n},a_{n})\;Q_{g_{\phi}}({\bm{x}}_{n},a_{n})\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ (45)
$=\sum_{{\bm{x}}_{n}\in\mathcal{X}_{n}}p_{g_{\phi}}({\bm{x}}_{n})\sum_{a_{n}\in\mathcal{A}_{n}}g_{\phi}(a_{n}\mid{\bm{x}}_{n})\;Q_{g_{\phi}}({\bm{x}}_{n},a_{n})\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ ($p({\bm{x}}_{n},a_{n})=p({\bm{x}}_{n})g(a_{n}\mid{\bm{x}}_{n})$)
$=\sum_{{\bm{x}}_{n}\in\mathcal{X}_{n}}p_{g_{\phi}}({\bm{x}}_{n})\sum_{a_{n}\in\mathcal{A}_{n}}Q_{g_{\phi}}({\bm{x}}_{n},a_{n})\;\nabla_{\phi}g_{\phi}(a_{n}\mid{\bm{x}}_{n}),$ ($g\,\nabla\log g=\nabla g$)
where $p_{g_{\phi}}({\bm{x}}_{n})=\Pr_{\tau\sim p_{g_{\phi}}(\cdot\mid{\bm{x}}_{L})}(X_{n}={\bm{x}}_{n})$ and $p_{g_{\phi}}({\bm{x}}_{n},a_{n})=p_{g_{\phi}}({\bm{x}}_{n})\,g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ are the marginals induced by the rollout distribution. Therefore,
$\nabla_{\phi}J(\phi)$ $=\mathbb{E}_{\tau\sim p_{g_{\phi}}(\cdot\mid{\bm{x}}_{L})}\!\left[r({\bm{x}}_{0})\sum_{n=1}^{L}\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\right]$
$=\sum_{{\bm{x}}\in\mathcal{X}_{1}\cup\mathcal{X}_{2}\dots\cup\mathcal{X}_{L}}p_{g_{\phi}}({\bm{x}})\sum_{a}Q_{g_{\phi}}({\bm{x}},a)\nabla_{\phi}g_{\phi}(a|{\bm{x}}),$
which ends the proof. ∎
This can also be proved in another way, just iteratively deriving the $\nabla_{\phi}V_{g_{\phi}}$. Using Lemma 2, we can prove our main objective as follows:
See 1
Proof.
We derive $\nabla_{\phi}\mathcal{L}_{\rm{output}}$ and $\nabla_{\phi}\mathcal{L}_{\rm{token}}$ separately. Then, by applying Lemma 2 to $\nabla_{\phi}\mathcal{L}_{\rm{output}}$, we obtain a form that matches $\nabla_{\phi}\mathcal{L}_{\rm{token}}$. First, expanding $\mathcal{L}_{\rm{output}}$ shows that it is equivalent to maximizing the advantage from the initial state:
$\mathcal{L}_{\rm{output}}(\phi)$ $=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{{\bm{x}}_{0}\sim p_{g_{\phi_{\rm{old}}}}(\cdot\mid{\mathbf{q}})}\bigr[\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0}\mid{\mathbf{q}})}A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (46)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0}\in\mathcal{X}}p_{g_{\phi_{\rm old}}}({\bm{x}}_{0}\mid{\mathbf{q}})\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0}\mid{\mathbf{q}})}A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (47)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0}\in\mathcal{X}}p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (48)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0}\in\mathcal{X}}\sum_{{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (49)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (50)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[A({\mathbf{q}},{\bm{x}}_{0})\bigl]$ (51)
Now we derive $\nabla_{\phi}\mathcal{L}_{\rm{output}}$ as follows:
$\nabla_{\phi}\mathcal{L}_{\rm{output}}(\phi)$ $=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[A({\mathbf{q}},{\bm{x}}_{0})\bigl]\biggl]$ (52)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})A({\mathbf{q}},{\bm{x}}_{0})\bigl]\biggl]$ (53)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\cdot\frac{(r({\mathbf{q}},{\bm{x}}_{0})-r_{g_{\phi_{\rm old}}}({\mathbf{q}}))}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\bigl]\biggl]$ (54)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\frac{r({\mathbf{q}},{\bm{x}}_{0})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\bigl]\biggl]$
$\hskip 56.9055pt-\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\frac{r_{g_{\phi_{\rm old}}}({\mathbf{q}}))}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\bigl]\biggl]$ (55)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\frac{r({\mathbf{q}},{\bm{x}}_{0})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\bigl]\biggl]$
$\hskip 56.9055pt-\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{r_{g_{\phi_{\rm old}}}({\mathbf{q}}))}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\nabla_{\phi}\bigr[\underbrace{\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}_{=1}\bigl]\biggl]$ (56)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\bigr[\sum_{{\bm{x}}_{0},{\bm{x}}_{1},...{\bm{x}}_{L}}p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\frac{r({\mathbf{q}},{\bm{x}}_{0})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\bigl]\biggl]$ (57)
$=\nabla_{\phi}\biggr[\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\frac{\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl]=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{\nabla_{\phi}V_{g_{\phi}}({\mathbf{q}})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl],$ (58)
where we define state-wise value function as the expectation of terminal reward starting from the state, i.e., $V_{g_{\phi}}({\bm{x}}_{n}|{\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:{n-1}}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}},{\bm{x}}_{n})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$ and we denote value function at the initial state as $V_{g_{\phi}}({\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$. Accordingly, state-wise action-value function can be defined as $Q_{g_{\phi}}({\bm{x}}_{n},a_{n}|{\mathbf{q}})=\sum_{{\bm{x}}_{n-1}}\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot V_{g_{\phi}}({\bm{x}}_{n-1}|{\mathbf{q}})$. Please note that Eq. (52)-57 is conceptually equivalent to reversing the derivation process of REINFORCE with baseline (Williams, 1992). Since we want to show that $\nabla_{\phi}\mathcal{L}_{\rm{token}}\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}\nabla_{\phi}\mathcal{L}_{\rm{output}}$, we breakdown $\nabla_{\phi}V_{g_{\phi}}({\mathbf{q}})$ using Lemma 2:
$\nabla_{\phi}\mathcal{L}_{\rm{output}}(\phi)$ $=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{\nabla_{\phi}V_{g_{\phi}}({\mathbf{q}})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl]$ (59)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{\sum_{\bm{x}}p_{g_{\phi}}({\bm{x}}|{\mathbf{q}})\sum Q_{g_{\phi}}({\bm{x}},a|{\mathbf{q}})\nabla_{\phi}g_{\phi}(a|{\bm{x}})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl]$ (60)
$\approx\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{\sum_{\bm{x}}p_{g_{\phi_{\rm{old}}}}({\bm{x}}|{\mathbf{q}})\sum Q_{g_{\phi_{\rm{old}}}}({\bm{x}},a|{\mathbf{q}})\nabla_{\phi}g_{\phi}(a|{\bm{x}})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl].$ (61)
where $\bm{x}$} comes from every state, i.e., ${\bm{x}}\in\mathcal{X}_{0}\cup\mathcal{X}_{1}\cup\dots\cup\mathcal{X}_{L}$. Here, Eq. (59) = Eq. (61) exactly holds at the first optimization step where $\phi=\phi_{\rm{old}}$. On the other side, we now show that $\nabla_{\phi}\mathcal{L}_{\rm{token}}$ is equivalent to Eq. (61). We start by converting the expectation into a probability-weighted sum as follows:
$\mathcal{L}_{\rm{token}}(\phi)=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\mathbb{E}_{{\bm{x}}_{0:L,a_{1:L}}\sim p_{g_{\phi_{\rm{old}}}}(\cdot\mid{\mathbf{q}})}\Biggl[\sum_{n=1}^{L}\frac{g_{\phi}(a_{n}\mid{\bm{x}}_{n})}{g_{\phi_{\rm old}}(a_{n}\mid{\bm{x}}_{n})}A({\mathbf{q}},{\bm{x}}_{0})\Biggr]$ (62)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\Biggl[\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\prod_{n=1}^{L}\biggl(\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot g_{\phi_{\rm{old}}}(a_{n},{\bm{x}}_{n})\biggr)\biggl(\sum_{n=1}^{L}\frac{g_{\phi}(a_{n}\mid{\bm{x}}_{n})}{g_{\phi_{\rm old}}(a_{n}\mid{\bm{x}}_{n})}A({\mathbf{q}},{\bm{x}}_{0})\biggr)\Biggr].$ (63)
For further explanation, we want to recap $p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{n-1}|{\bm{x}}_{n})=\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})$. Please note that $p_{g_{\phi_{\rm{old}}}}>0$ only if ${\bm{x}}_{n}^{a_{n}}=\mathbb{M}$ and ${\bm{x}}_{n-1}^{a_{n}}\neq\mathbb{M}$ since $g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})=0$ for ${\bm{x}}_{n}^{a_{n}}\neq\mathbb{M}$, $\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})=0$ for all ${\bm{x}}_{n-1}$ satisfying $x_{n-1}^{a_{n}}=\mathbb{M}$ by definition. To simplify the derivation process, we consider $\widehat{\mathcal{L}}_{\rm{token,i}}(\phi,{\mathbf{q}})$ defined as follows:
$\widehat{\mathcal{L}}_{\rm{token,i}}(\phi,{\mathbf{q}})=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}|{\mathbf{q}})\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}r({\mathbf{q}},{\bm{x}}_{0})$ (64)
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigl(\prod_{j=1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigr)\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (65)
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (66)
$\hskip 28.45274pt\cdot\Bigr(p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{i-1}|{\bm{x}}_{i},{\mathbf{q}})\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (67)
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (68)
$\hskip 28.45274pt\Bigr(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i})\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (69)
$=\sum_{\begin{subarray}{c}{\bm{x}}_{i},\dots,{\bm{x}}_{L}\\ a_{i+1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (70)
$\cdot\sum_{\begin{subarray}{c}{\bm{x}}_{i-1},\dots,{\bm{x}}_{0}\\ a_{i},\dots,a_{1}\end{subarray}}\biggl(\Bigr(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i})\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\biggr)\Biggr)$ (71)
$=\sum_{\begin{subarray}{c}{\bm{x}}_{i},\dots,{\bm{x}}_{L}\\ a_{i+1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (72)
$\cdot\sum_{a_{i}}\sum_{{\bm{x}}_{i-1}}\biggl(g_{\phi}(a_{i}\mid{\bm{x}}_{i})\cdot\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})\sum_{\begin{subarray}{c}{\bm{x}}_{i-2},\dots,{\bm{x}}_{0}\\ a_{i-1},\dots,a_{1}\end{subarray}}\biggr(\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\biggr)\Biggr)$ (73)
$=\sum_{{\bm{x}}_{i},...{\bm{x}}_{L}}\sum_{a_{i+1},...,a_{L}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (74)
$\cdot\sum_{a_{i}}g_{\phi}(a_{i}\mid{\bm{x}}_{i})\cdot\sum_{{\bm{x}}_{i-1}}\biggl(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})\sum_{\begin{subarray}{c}{\bm{x}}_{i-2},\dots,{\bm{x}}_{0}\\ a_{i-1},\dots,a_{1}\end{subarray}}\biggr(\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\biggr)\Biggr)$ (75)
$=\sum_{{\bm{x}}_{i},...{\bm{x}}_{L}}\sum_{a_{i+1},...,a_{L}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (76)
$\hskip 113.81102pt\cdot\sum_{a_{i}}g_{\phi}(a_{i}\mid{\bm{x}}_{i})\cdot\sum_{{\bm{x}}_{i-1}}\biggl(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})\cdot V_{\phi_{\rm{old}}}({\bm{x}}_{i-1}|{\mathbf{q}})\biggr)\Biggr)$ (77)
$=\mathbb{E}_{{\bm{x}}_{i}\sim p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})}[\sum g_{\phi}(a_{i}|{\bm{x}}_{i})\cdot Q_{g_{\phi_{\rm{old}}}}({\bm{x}}_{i},a_{i}|{\mathbf{q}})]$ (78)
Then,
$\nabla_{\phi}\mathcal{L}_{\rm{token}}(\phi)$ $=\sum_{i=1}^{L}\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\Biggl[\frac{\widehat{\mathcal{L}}_{\rm{token,i}}(\phi)}{{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}}\Biggr]$ (79)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\frac{\sum_{i=1}^{L}\mathbb{E}_{{\bm{x}}_{i}\sim p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})}[\sum\nabla_{\phi}g_{\phi}(a_{i}|{\bm{x}}_{i})\cdot Q_{g_{\phi_{\rm{old}}}}({\bm{x}}_{i},a_{i}|{\mathbf{q}})]}{{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}}$ (80)
$=\mathbb{E}_{{\mathbf{q}}\sim\rho_{\mathcal{Q}}}\biggr[\frac{\sum_{\bm{x}}p_{g_{\phi_{\rm{old}}}}({\bm{x}}|{\mathbf{q}})\sum Q_{g_{\phi_{\rm{old}}}}({\bm{x}},a|{\mathbf{q}})\nabla_{\phi}g_{\phi}(a|{\bm{x}},{\mathbf{q}})}{\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon}\biggl].$ (81)
Eq. (79) holds where $r_{g_{\phi_{\rm{old}}}}({\mathbf{q}})$ in $A({\mathbf{q}},{\bm{x}}_{0})=(r(q,{\bm{x}}_{0})-r_{g_{\phi_{\rm{old}}}}({\mathbf{q}}))/(\mathrm{std}_{g_{\phi_{\rm old}}}({\mathbf{q}})+\epsilon)$ can be deleted as we have done in $\nabla_{\phi}\mathcal{L}_{\rm{output}}$. Since Eq. (61) and Eq. (81) are identical, we conclude $\nabla_{\phi}\mathcal{L}_{\rm{token}}\approx\nabla_{\phi}\mathcal{L}_{\rm{output}}$ and $\nabla_{\phi}\mathcal{L}_{\rm{token}}=\nabla_{\phi}\mathcal{L}_{\rm{output}}$ at the first optimization step where $\phi=\phi_{\rm{old}}$. ∎
C.3 Surrogate loss of output-level KL divergence
Our final goal is to provide evidence that by performing gradient ascent on $\mathcal{L}_{\rm{KL}}$, we can expect to decrease $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$. We provide two propositions here. The first proposition shows that trajectory-wise KL divergence is the upper bound of output-level KL divergence. The second proposition shows that the gradient of trajectory-wise KL divergence is equivalent to $\mathcal{L}_{\rm{KL}}$.
Proposition 2.
For MDMs where $p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})=\sum_{{\bm{x}}_{1},...,{\bm{x}}_{L}}p_{\phi}({\bm{x}}_{0:L}|{\mathbf{q}})$, consider the terminal KL divergence $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$ and trajectory-wise KL divergence $D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right)$ defined as follows:
$D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)=\mathbb{E}_{\tau\sim p_{g_{\rm{\phi}}}({\bm{x}}_{0}|{\mathbf{q}})}\biggl[\log\frac{p_{g_{\rm{\phi}}}({\bm{x}}_{0}|{\mathbf{q}})}{p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})}\biggr],$ (82)
$D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right)=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}\biggl[\log\frac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})}\biggr]$ (83)
Then, $D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right)$ is upper bound of $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$:
$D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)<D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right).$ (84)
Proof.
$D_{\mathrm{KL}}$ $\left(p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})\right)=\mathbb{E}_{{\bm{x}}_{0}\sim p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}\left[\log\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})}\right]$ (85)
$=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\left[\log\frac{p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})}\right]$ (86)
$=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\left[\log\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}-\log\frac{p_{g_{\phi}}({\bm{x}}_{1:L}\mid{\bm{x}}_{0},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{1:L}\mid{\bm{x}}_{0},{\mathbf{q}})}\right]$ (87)
$=D_{\mathrm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}}),|,p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)$ (88)
$-\mathbb{E}_{{\bm{x}}_{0}\sim p_{g_{\phi}}({\bm{x}}_{0}\mid{\mathbf{q}})}\left[D_{\mathrm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{1:L}\mid{\bm{x}}_{0},{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{1:L}\mid{\bm{x}}_{0},{\mathbf{q}})\right)\right]$ (89)
$\leq D_{\mathrm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right).$ (90)
Therefore, we can lower $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$ by decreasing $D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right)$. Now we show that $\nabla_{\phi}D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}|{\mathbf{q}})\right)=\nabla_{\phi}\mathcal{L}_{\rm{KL}}$:
Proposition 3.
Consider the trajectory-wise KL divergence $D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)$ and tractable $\mathcal{L}_{\rm{KL}}$ defined as follows:
$\mathcal{L}_{\rm KL}({\bm{x}}_{0:L},a_{1:L},{\mathbf{q}})={\rm StopGrad}\!\Bigl(\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\phi_{\rm old}}}({\bm{x}}_{0:L}|{\mathbf{q}})}\cdot\bigl(1+\log\tfrac{p_{g_{\phi}}({\bm{x}}_{0:L}|{\mathbf{q}})}{p_{g_{\rm ref}}({\bm{x}}_{0:L}|{\mathbf{q}})}\bigr)\Bigr)\cdot\sum_{n=1}^{L}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})$ (91)
Then, the gradient of $D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)$ and $\mathbb{E}[\mathcal{L}_{\rm{KL}}]$ are equal:
$\nabla_{\phi}D_{\rm{KL}}\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)=\nabla_{\phi}\mathbb{E}_{{\bm{x}}_{0:L},a_{0:L}\sim p_{g_{\phi_{\rm{old}}}}}[\mathcal{L}_{\rm{KL}}]$ (92)
Proof.
$D_{\rm{KL}}$ $\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)$ (93)
$=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\!\left[\log\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\right]$ (94)
$=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\!\left[\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot\log\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\right]$ (95)
$=\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\!\left[\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\right]$ (96)
$\nabla_{\phi}D_{\rm{KL}}$ $\left(p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})||p_{g_{\mathrm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})\right)$ (97)
$=\nabla_{\phi}\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\!\left[\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\right]$ (98)
$=\mathbb{E}\Biggl[\nabla_{\phi}\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}$ (99)
$\hskip 113.81102pt+\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot\nabla_{\phi}\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\Biggr]$ (100)
$=\mathbb{E}\Biggl[\biggr(\sum_{i=1}^{L}\frac{\nabla_{\phi}p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\biggl)\biggr(\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\biggl)\cdot\biggl(\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\biggr)$ (101)
$\hskip 142.26378pt+\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\biggr(\sum_{i=1}^{L}\frac{\nabla_{\phi}p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\biggl)\Biggr]$ (102)
$=\mathbb{E}\Biggl[\biggr(\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\biggl)\cdot\biggl(1+\sum_{i=1}^{L}\log\frac{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\mathrm{ref}}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\biggr)\biggr(\sum_{i=1}^{L}\frac{\nabla_{\phi}p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}{p_{g_{\phi}}({\bm{x}}_{i-1}\mid{\bm{x}}_{i},{\mathbf{q}})}\biggl)\Biggr]$ (103)
$=\nabla_{\phi}\mathbb{E}\Biggl[{\rm{StopGrad}}\Bigl(\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}\cdot(1+\log\frac{p_{g_{\phi}}({\bm{x}}_{0:L}\mid{\mathbf{q}})}{p_{g_{\rm{ref}}}({\bm{x}}_{0:L}\mid{\mathbf{q}})})\Bigr)\cdot\sum_{i=1}^{L}\log g_{\phi}(a_{i}|{\bm{x}}_{i})\Biggr]$ (104)
$=\nabla_{\phi}\mathbb{E}[\mathcal{L}_{\rm{KL}}]$ (105)